OpenStudy (anonymous):

simplify 48w^10x^8/18w^3x^2

OpenStudy (anonymous):

\[48w ^{10} x^8 / 18x^3x^2\] 48/18 = 8/3 10 - 3 = 7 8 - 2 = 4 so \[8x^7 y^4\]

OpenStudy (anonymous):

wouldn't the y be 6th power

OpenStudy (anonymous):

Yes it would be, sorry about that,

OpenStudy (anonymous):

then would you use the 3 as a denominator and not have w and x?

OpenStudy (anonymous):

Yes I think so, let me rewrite what I typed up there cuz its confusing

OpenStudy (anonymous):

8w^7x^6/3 ? I am sorry to be so dense.

OpenStudy (anonymous):

\[8w^7 y^6 /3\] Being that way is good! Trust me, it decreases the chances of getting a wrong answer by a ton!

OpenStudy (anonymous):

thank you so very much.

OpenStudy (anonymous):

You're really good at this:) I think you helped me more lol:D no problem!

OpenStudy (anonymous):

no i am not